OpenStudy (anonymous):

Are there better approximations of a function at some value of x besides a local linearization?

OpenStudy (ikram002p):

yeah there is like many approximation functions/ series

OpenStudy (anonymous):

Are the Fourier series one of them?

OpenStudy (ikram002p):

yep :) also tylor series , geometric series mmm also there is newtons method / bisector method and other cool stuff

OpenStudy (ikram002p):

search for Numerical analysis ull gonna find very intresteing algorithms for approximation functions

OpenStudy (anonymous):

Nice to know, I only know the Fourier series because of exploring stuff, I'm only inCalc BC I haven't gotten to the power series and the taylor series I always keep hearing about and seeing, but not understanding! -_-

OpenStudy (ikram002p):

ic , well check Numerical analysis , the algorithms there are soo easy u dnt even need to know calcules to solve :)

OpenStudy (anonymous):

When I was in pre-calc, I wanted to understand where they got \[\Large e^{i \theta}={\cos(\theta)+isin(\theta)}\] from, but I always got a calculus proof with the taylor series and all that and I'm like woah, until I found a page that broke it down, w/o calculus

OpenStudy (ikram002p):

haha , yeah this is in complex analysis =")